OpenStudy (anonymous):

on a meter stick a fulcrum is placed at 50cm mark. Because of this fulcrum the meter stick can rotate about it. Then a 3kg mass is hung from the meter stick at 10cm mark. Where should a 5kg mass be placed on the meter stick so that the meter stick is in equilibrium?

OpenStudy (anonymous):

 the system is in rotational equilibrium about the fulcrum so, \[\Sigma\tau=0\]

OpenStudy (anonymous):

i understand it at equilibrium but what is the location on the meter stick of the 2nd mass??

OpenStudy (anonymous):

find "x"

OpenStudy (anonymous):

set up the equation first.. clockwise torques negative and counter cllockwise positive

OpenStudy (anonymous):

\[\Sigma=(3)(10)-(5)(x)=0\]

OpenStudy (anonymous):

nope... distances from the point of rotation... 3(50-10)-(5)(x-50)=0

OpenStudy (anonymous):

okay got it the answer is 74cm??

OpenStudy (anonymous):

sounds reasonable.. x-50 should be less than (50-10)

OpenStudy (anonymous):

got it thanks!

OpenStudy (anonymous):

a mass 20kg starts from rest at the top of the an incline. This mass reaches the bottom of the incline(30 degrees) with a speed of 9.90 m/s. What is the height of the incline ? What is the length of the incline? friction less surface. I got 5m as the height how do i get the length?